using base 2 logarithms solve the following equation
x-3
5 =32

Answer by stanbon(75887)

using base 2 logarithms solve the following equation
x-3
5 =32
(x-3)log2(5) = log2(32)
---------------------------
(x-3)log2(5) = 5
x-3 = 5/log2(5)
----
x = [5/log2(5)] + 3
----------------------
Cheers,
Stan H.